# The Unseen Frontier: Pushing the Limits of LLM Sparsity with Surrogate-Free ADMM

### v1

@@ B.1 Implementation and reproduction details
All experimental results in this work are obtained with unified codebase, while baseline methods are reproduced using their original implementations whenever available. The environment configuration (dependencies, versions, and training scripts) will be made available at https://github.com/log-postech/elsa.
@@
To obtain baseline results (Wanda, SparseGPT, ALPS, L-ADMM, SAFE, SparseLLM), we follow the convention of Frantar & Alistarh (2023), sampling 128 calibration sequences from the C4 dataset with sequence length 2048. For Elsa, we adopt the same strategy, but use larger calibration sets to account for the iterative nature of our optimization.
@@
We train Elsa for 4,096 steps with a batch size of 8 across all model scales, using Adam as the base optimizer. The penalty parameter follows a cosine schedule, gradually increasing from 0 at the start to $\lambda$ at the end of training. All model parameters and optimizer states uses full precision for training (except for memory-efficient setting and ablations), and automatic mixed precision with bf16 precision is used for efficient training. A full list of hyperparameter configurations is provided in Tables˜4 and 5.
@@
Perplexity is measured on the held-out (validation) C4 (Raffel et al., 2020) and Wikitext2 (Merity et al., 2017) datasets. Zero-shot performance is evaluated with lm-eval-harness across seven standard tasks: ARC-Easy/Challenge (ARC-E/C) (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OpenBookQA (OBQA) (Mihaylov et al., 2018), RTE (Zeng & Urtasun, 2018), and Winogrande (Sakaguchi et al., 2021), and we report the average accuracy as in Section˜5.1.
@@ Appendix C Additional results
Here we provide numerical results used to make visual plots in the main text, and additional result reporting LLaMA-2-13B zero-shot task accuracy.

### v2

@@ B.1 Implementation and reproduction details
All experimental results in this work are obtained with unified codebase, while baseline methods are reproduced using their original implementations whenever available. The environment configuration (dependencies, versions, and training scripts) can be found at https://github.com/log-postech/elsa.
@@
To obtain baseline results (Wanda, SparseGPT, ALPS, L-ADMM, SAFE, SparseLLM), we follow the convention of Frantar and Alistarh (2023), sampling 128 calibration sequences from the C4 dataset with sequence length 2048. For Elsa, we adopt the same strategy, but use larger calibration sets to account for the iterative nature of our optimization.
@@
We train Elsa with 32,768 data points, where each data point has sequence length of 2048. Batch size and number of steps differ across model, which is presented at table Table˜6. We use Adam as the base optimizer. The penalty parameter is kept constant for moderate sparsity levels (50-60%), while we use cosine schedule, which gradually increases the penalty parameter from 0 at the start to $\lambda$ at the end of training. All model parameters and optimizer states uses full precision for training (except for memory-efficient setting and ablations), and automatic mixed precision with bf16 precision is used for efficient training. A full list of hyperparameter configurations is provided in Tables˜4 and 5.
@@
Perplexity is measured on the held-out (validation) C4 (Raffel et al., 2020) and Wikitext2 (Merity et al., 2017) datasets. Zero-shot performance is evaluated with lm-eval-harness across seven standard tasks: ARC-Easy/Challenge (ARC-E/C) (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OpenBookQA (OBQA) (Mihaylov et al., 2018), RTE (Zeng and Urtasun, 2018), and Winogrande (Sakaguchi et al., 2021), and we report the average accuracy as in Section˜5.1.
@@
For the baseline (Wanda+LoRA / Full), we first prune the pretrained model with Wanda, and then retrain the remaining parameters using either LoRA (Hu et al., 2022) or full fine-tuning (Full). For a fair comparison, we use the same number data points with same sequence length as Elsa, with hyperparameters tuned separately for each method.
B.4 Details for Section˜5.4
@@ Appendix C Additional results
Here we provide additional results for different sparsity patterns (Section˜C.1), ablation studies on objective function and projection step (Section˜C.2), and a numerical results used to make visual plots in the main text followed by additional result reporting LLaMA-2-13B zero-shot task accuracy.
Method Wiki()($\downarrow$) C4()($\downarrow$)
SparseGPT 85.84 98.35
OWL 48.07 52.32
EvoPress 28.76 33.72
Elsa (EvoPress) 26.11 29.33
Elsa 24.97 29.09
C.1 Other sparsity patterns
Perplexity ($\downarrow$) Tasks ($\uparrow$)
0% Dense 5.47 7.26 43.35 76.26 77.68 57.14 31.40 62.82 69.06 59.67
Wanda 12.13 15.63 30.46 61.83 68.26 41.28 24.20 53.07 62.51 48.80
SparseGPT 10.87 13.61 30.97 64.06 67.61 43.47 24.20 56.32 66.38 50.43
L-ADMM 10.19 12.51 32.85 66.04 68.81 45.05 25.40 56.32 66.38 51.55
ALPS 9.945 12.09 34.47 68.86 73.79 49.40 27.60 55.60 67.25 53.85
SAFE 9.914 12.53 30.46 63.43 66.42 44.66 21.60 53.07 61.80 48.78
SparseLLM 11.29 13.95 30.55 61.91 71.10 43.62 24.40 57.40 65.82 50.69
Elsa 10.15 12.34 31.49 61.24 66.36 47.87 23.60 52.71 63.85 49.59
Wanda 8.603 11.33 34.47 67.05 72.87 46.98 26.80 54.15 66.93 52.75
SparseGPT 8.508 10.81 34.81 68.56 71.77 48.26 27.80 56.68 68.11 53.71
L-ADMM 8.12 10.37 35.58 68.18 72.48 49.45 28.80 58.12 67.17 54.25
ALPS 8.103 10.29 33.28 65.19 68.75 45.96 26.20 55.96 65.98 51.62
SAFE 8.043 10.47 31.57 66.84 68.04 48.55 23.40 53.07 65.04 50.93
SparseLLM 8.679 11.04 34.90 68.35 75.14 48.28 26.20 56.68 66.46 53.71
Elsa 9.20 11.47 32.25 64.69 69.42 49.90 27.40 53.07 63.22 51.42
We first evaluate Elsa for its adaptivity to N:M semi-structured sparsity, a setting designed for some current hardwares to accelerate computations (Sun et al., 2024; Fang et al., 2024). The results of both perplexity and zero-shot prediction accuracy are reported in Table˜8. Elsa is roughly on par with existing methods, and yet, it is noteworthy that these 2:4 and 4:8 sparsity patterns only ensure 50% sparsity. More importantly, these results indicate that Elsa can easily adapt to arbitrary constraints of moderate sparsity levels without much trouble.
We also compare Elsa with non-uniform sparsity allocation based pruning methods. Specifically, we compare to OWL (Yin et al., 2024a) that allocates sparsity based on outlier distributions and to EvoPress (Sieberling et al., 2024) that uses an evolutionary search strategy to determine the non-uniform sparsity levels over different layers. We further set up a method that overrides Elsa with the mask found by the evolutionary strategy of EvoPress. Note that the sparsity level is set to be 70%; it is simply because these methods only works or reports up to this level. The results are presented in Table˜7. One can see that Elsa substantially outperforms OWL and shows an improvement over EvoPress as well: to elaborate, for instance, it achieves the C4 perplexity of $29.09$, compared to $33.72$ for EvoPress and $52.32$ for OWL. Notably, adopting the non-uniform mask found by EvoPress within Elsa yields some gains over the EvoPress itself, but it still falls short of the uniform allocation in Elsa, demonstrating the strength of our surrogate-free global formulation.
C.2 Ablations
Figure 6: Effect of NTP on data efficiency and perplexity.
Table 9: Effectiveness of geometric projection (✓).
Sparsity
70% 29.44 28.24
80% 40.06 37.50
90% 65.41 48.69
We also evaluate the effectiveness of the objective-aware projection on high-sparsity regimes. Specifically, we measure the perplexity of LLaMA-3.2-3B model pruned for 70-90% sparsity levels by turning on and off of the projection and report the results in Table˜9. The benefit of objective-aware projection grows with sparsity: perplexity gap increases from 1.20 at 70% sparsity to 2.56 at 80%, and widens further at 90%. This demonstrates that incorporating objective-aware importance into the projection step can be beneficial particularly in high sparsity regimes.
C.3 Numerical results
Here we provide numerical results used to produce visual figures on the main text, followed by additional results reporting LLaMA-2-7B/13B zero-shot task accuracy.